When a lightbulb is turned on, 6.70 C of charge pass through the filament of the bulb in 10.0 s. How many electrons pass through the filament in 5 s?
Answer #1

I=Q/t=6.70/10=0.670 A

Now,0.670=q/5

Q=.670 ×5=3.35 Column.

Note:- Q - charge,,i - current,,t- time